A $100,000 dollar face value bond pays 4% interest annually for the next 30 years. The market rate is also 4%,If the bond is strippedWhat will the lump sum and Annuity Sell for?
Answer #1

Since the interest rate on bond and market rate are same, the bond will sell at par value i.e. $100,000

Value of lump sum = Face Value*Present value factor

= 100,000*PVF(4%, 30 years)

= 100,000*0.308319

= $30,831.9

Hence, value of lump sum = $30,831.9

Value of Annuity = 100,000 - 30,831.9 = $69,168.1